what are the 2 consecutive numbers that equal 205
Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Only 205 is equal to 205.
----------
If you mean 2 consecutive integers whose sum = 205:
205/2 = 102.5
--> 102 & 103
----------------
But, you didn't specify integers, either, which injects more confusion.
-------------
It's important to know what you want to do.
Not knowing what you're trying to do reduces your chances of success.

Answer by ewatrrr(24785)


Hi,
*Note: consecutive numbers differ by 1
Let n be the first, then (n + 1) would be the other
the question states
n + (n + 1) = 205
2n + 1 = 205
2n = 204
n = 102
The 2 consecutive numbers would be 102 and 103
check your answer
102 + 103 = 205